OpenStudy (anonymous):

Solve the system of equations and enter the solution as an ordered pair. 4x - 4y + 7 = 15 8y = 2x + 14

OpenStudy (anonymous):

solve for y in the second equation : y=1/4x + 7/4

OpenStudy (anonymous):

then substitute this into the top equation 4x-4(1/4x+7/4)=15

OpenStudy (anonymous):

then substitute this into the top equation 4x-4(1/4x+7/4)=15